OpenStudy (anonymous):

The sun’s mass is about 2.7 x 107 times greater than the moon’s mass. The sun is about 400 times farther from Earth than the moon. How does the gravitational force exerted on Earth by the sun compare with the gravitational force exerted on Earth?

OpenStudy (ookawaiioo):

F1 = G(M1*M2/r^2) F2 = G(M1*M3/r^2) M1 = earth M2 = sun M3 = moon F1/F2 = the ratio of the two

OpenStudy (ookawaiioo):

G will cancel out

OpenStudy (anonymous):

Im sorry im still a bit confused. I am not sure how to proceed @oOkawaiiOo

OpenStudy (ookawaiioo):

G(M1*M2/r^2) /G(M1*M3/r^2) G and r^2 will cancel out. Solve from there